Please help me in figuring out this problem: If N=94, find the temperature in degrees Fahrenheit, T.
T=Temp. in degrees Fahrenheit
N= number of chirps per minute
T=50+ N-40/4
Thank you,

Answer by ewatrrr(24785)

T = 50+ 54/4 = 50 + 13.5 = 63.5 degrees Fahrenheit